OpenStudy (anonymous):

x+y-z=5 2x-y+3z=-3 2x+3y=-4

OpenStudy (anonymous):

could you help me out, kelligirl?

OpenStudy (anonymous):

you writing down the problem to go over it step by step with me?

OpenStudy (anonymous):

i think she is writing the whole solution!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x + y - z = 5 ---->(3) 2x - y + 3z = -3 -------------- 3x + 3y - 3z = 15 (result of multiplying by 3) 2x - y + 3z = -3 --------------add 5x + 2y = 12 2x + 3y = -4 -->(-2) 5x + 2y = 12 -->(3) ---------------- -4x - 6y = 8 (result of multiplying by -2) 15x + 6y = 36 (result of multiplying by 3) -----------------add 11x = 44 x = 44/11 x = 4 2x + 3y = -4 2(4) + 3y = -4 8 + 3y = -4 3y = -4 - 8 3y = -12 y = -12/3 y = -4 x + y - z = 5 4 - 4 - z = 5 -z = 5 z = -5 check... x + y - z = 5 4 - 4 -(-5) = 5 0 + 5 = 5 5 = 5 (correct) x = 4, y = -4, z = -5

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but

OpenStudy (anonymous):

ehh

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

if you have any questions, please ask

OpenStudy (anonymous):

oh yes

OpenStudy (anonymous):

I should have said: solve the equation using Cramer's rule

OpenStudy (anonymous):

like this

OpenStudy (anonymous):